# Anisotropic Spin Ice on a Breathing Pyrochlore Lattice

### v1

@@ V.1 Monte Carlo sampling
We collect states by Monte Carlo sampling a system of size $L=10$ (corresponding to $N_{\text{spins}}=16000$ spins) and compute the neutron scattering response. For $T\rightarrow 0$ states can be directly computed by sampling within the ground state manifold. To this end, sampling algorithms that allow us to generate all possible ground states of an anisotropic phase are employed. These sampling algorithms have a transition probability of $1$ when starting from a state within the ground state manifold and are described in detail in App. E.
@@
V.2 Mean field analysis
@@
Here, $L$ is the linear system size, and we insert the Fourier transform of our interaction Hamiltonian, specified in App. C. The Lagrange multiplier $\lambda$ can be interpreted as the thermal occupation value of the lowest energy mode [56] and will drop to zero if there is a phase transition in the large-$N$ limit [56, 54]. An analysis of the Lagrange multiplier in the different $\delta$-regimes can be found in App. C and is consistent with the findings of section IV.
The scattering result in the SF channel for $N\rightarrow\infty$ is computed by inserting the correlator Eq. (8) into Eq. (6). To resolve this at different temperatures, we first need to solve for $\lambda$ in (9) at the necessary inverse temperature $\beta$. We can also compare the result from the large-N approximation to numerical data, where we sample ground states for each system according to the ground state rule. In the mean-field approximation, we compute the correlations at a temperature well below the crossover ($T\ll 1J$) to accurately reproduce correlations close to or at the ground state.
@@ V.3 Results
The normalized SF channel of the neutron scattering cross section for the six different phases at or close to the ground state is shown in Fig. 4. The intensity is normalized for each system, $(S_{\text{SF}}(\mathbf{q})-\text{min}(S_{\text{SF}}(\mathbf{q})))/(\text{max}(S_{\text{SF}}(\mathbf{q}))-\text{min}(S_{\text{SF}}(\mathbf{q})))$, such that the signatures will lie between 00 and $1$. We can see that for any limit of $\delta_{A},\delta_{B}$, the spin structure factor computed numerically shows no significant difference from the result obtained by the SCGA when the temperature is set close to the transition temperature. It is important to note that the signatures on the antidiagonal in the [hl0] cut are due to the neutron polarization being almost parallel to the scattering wave vector and not a signature of the underlying spin configuration.
Figure 4: Normalized spin structure factor in Spin-Flip (SF) channel. Columns show the same system, and rows have the same cut in the Brillouin zone. Colums: 1 Isotropic, 2 Plane Ice, 3 Omni-Plane, 4 Line-Order, 5 Plane-Paramagnetic, 6 Ferromagnetic. Rows: 1 [hhl], 2 [hl0], 3 [h0l]. Every panel shows Monte Carlo (MC) data on the top and mean field theory (MFT) on the bottom. The Monte Carlo data is obtained by sampling in the degenerate ground-state sector for systems with $L_{x}=L_{y}=L_{z}=10$ fcc unit cells, corresponding to $N_{\text{spin}}=16000$. The intensities are normalized across a single system between 00 and $1$ in a.u..
@@
In the planar ice phase, the pinch points with bow-tie patterns are no longer present in the [hhl] and [h0l] plane. Only in the [hl0] cut do pinch points still exist with a bow-tie pattern. This is understood from the cut being exactly parallel to the ice plane. Here, in the plane, correlations are still of dipolar form, with $C_{\alpha\beta}(\mathbf{r})\propto\frac{1}{r^{4}}(r^{2}\delta_{\alpha\beta}-3r_{\alpha}r_{\beta})$, for $r$ in the [hl0] plane. This will lead to a structure factor of the form $S_{\alpha\beta}\propto(\delta_{\alpha\beta}-\frac{k_{\alpha}k_{\beta}}{k^{2}})$ [62] in the plane, due to the square ice constraint in the ground state. The location of the pinch points in square ice are located at the high symmetry points of the square lattice, (±$(\pm 1,0)$, $(0,\pm 1)$, and (±$(\pm 1,\pm 1)$. This translates to pinch points in the [hl0] cut at (±$(\pm 100)$, $(0\pm 10)$, and (±$(\pm 1\pm 10)$ and symmetry equivalent points in the extended Brilloin zone. Since single planes are not correlated, other cuts in the lattice do not show this behavior.
@@
In the omni-plane case, spins order in particular patterns within the different tetrahedra-planes of the lattice. A single $A$ tetrahedra has a local magnetization vector that is either parallel or antiparallel to the $z$ axis. These tetrahedra will align predominately antiferromagnetically along planes perpendicular to the $z$ direction and predominantly ferromagnetically for planes perpendicular to the $x$ and $y$ direction.
@@
In a unit cell, there are six possible spin configurations. Two of which are fully ferromagnetic, leading to sharp Bragg peaks in the Brillouin zone center at $(000)$. The other four have all the $A$ tetrahedra in either the $x,z$ or the $y,z$-plane ferromagnetically aligned, while in all other planes, the $A$ tetrahedra are antiferromagnetically aligned. For these four states, however, two adjacent ferromagnetic layers have opposite polarization. This is necessary to get the antiferromagnetic alignment in the other layers. When we now project the two ferromagnetic layers of a unit cell down onto the plane, the two oppositely polarized planes form an antiferromagnetic square lattice. For this order, characteristic Bragg peaks at the four corners of the square lattice can be observed. On the other hand, projecting the antiferromagnetic layers of a unit cell onto the plane gives a square lattice with stripe order. For a square lattice with horizontal or vertical stripe order, Bragg peak at $(0,\pm 1)$ or (±$(\pm 1,0)$ appear.
@@
Any plane flip in the system will introduce line defects in the square lattice and thereby broaden the sharp Bragg peaks along the flipping direction. This explains the square grid structure in the [h0l] plane; the AFM square lattice peaks are broadened by introducing line defects. In the [hl0] plane, on the other hand, the stripe order peaks at (±$(\pm 1\pm 10)$ are broadened due to the introduction of line defects and form the cross-patterns.
@@
In the plane-paramagnetic case, spins form decoupled planes, where within planes, the local magnetization vector of the $B$ tetrahedra align antiferromagnetically. This order in the $x,y$-planes is visible as Bragg peaks in the [hl0] cut, where the $B$ tetrahedra form a square lattice. In a square lattice with AFM order, we expect Bragg-peaks to form at (±$(\pm 1,\pm 1)$; however, the square lattice formed by the $B$ tetrahedra of a unit cell is rotated by $\pi/2$ compared to the $x,y$-projection of the fcc unit cell. This leads to the peaks at the expected locations of (±$(\pm 100)$ and $(0\pm 10)$ as well as symmetry equivalent positions in the extended Brillouin zone. The uncorrelated planes can be inferred by the vertical lines in the [h0l] cut.
@@
Figure 5: Large-N spin structure factor in Spin-Flip (SF) channel at different temperatures. While for high temperatures ($T\gg 1$) the signatures for the isotropic (left) and anisotropic (right) systems agree, when the temperature is lowered (top to bottom), the features look significantly different. The line-order regime is evaluated with $\delta_{A}=\delta_{B}=0.05$ and the [hhl] cut in the Brillouin zone is shown. The scale is chosen uniformly across all temperatures and both systems.
While these signatures will be visible at low temperatures, very close to the ground state, remnants of the change in correlations will also be visible at higher temperatures. We show how the neutron scattering response resulting from the large-N approximation changes between isotropic spin ice and the anisotropic systems for decreasing temperatures with the example of the line-order case in Fig. 5. Here, the colorscale is chosen uniformly across all temperatures and both systems for better comparison. We can observe that at high temperatures, for $T\gg 1J$, the scattering data will be indistinguishable between the isotropic and the anisotropic cases. However, significant differences can be observed below the spin-ice crossover, where the anisotropic systems gradually enter their respective ground-state manifold and the aforementioned signatures become visible. Measuring the structure factor as a function of temperature thus directly unveils the anisotropic couplings of model Eq. (1).
@@ VI Summary and Outlook
In this work, we explored how anisotropic perturbations in the nearest-neighbor spin ice Hamiltonian affect the ground state structure and equilibrium properties of the spin system on the breathing pyrochlore lattice. By tuning the strain parameter, $\delta_{A/B}$, independently on the two tetrahedral sublattices, we demonstrated that the ground state degeneracy exhibits a sign-dependent change, leading to additional five distinct phases. In all cases, the degeneracy is reduced; strikingly, in only one of the five anisotropic phases, the degeneracy grows exponentially with the full system volume, while in the other scenarios, the ground state degeneracy scales with a lower-dimensional cut of the system.
@@
We also investigated the equilibrium properties of the system via the specific heat and entropy. Our results show that the anisotropic perturbation will either lead to a phase transition into a symmetry-broken phase at a critical temperature (omni-plane, plane paramagnetic, and ferromagnetic phase), a phase transition at zero temperature (line-order) or a crossover into a phase with decoupled planes (planar ice). In all cases but one, the planar ice phase, the residual entropy density approaches zero for increasing system sizes. The static spin structure factors of the anisotropic systems display significant deviations from the conventional spin ice case, showing only dipolar correlations in the planar ice phase.
Our analysis introduces a new perspective on frustrated magnetic systems, particularly in the context of strain engineering. While strain effects in spin ice systems have been previously studied — for instance, in Refs. [20, 19, 30, 17] — our introduction of sublattice-dependent anisotropy substantially enriches the field and opens avenues for further exploration.
@@
Looking forward, several intriguing questions arise from our findings. Since the ground state degeneracy is significantly reduced in some cases, an important question concerns the mechanisms by which equilibrium is achieved. Furthermore, the timescales on which equilibrium is approached are to be investigated. Additionally, in the ground state limit, certain lower-dimensional structures within the lattice (such as planes or lines) effectively decouple. This raises the question of how these decoupled structures influence the mobility of excitations in the system and their possible relations to “fractons” [63, 64, 65, 66, 67, 24, 25]?
Additionally, in this study, we limited our investigations to the nearest neighbor Ising Hamiltonian on the pyrochlore lattice and thereby neglected the long-range dipolar interactions between spins that contribute to the interactions in the typically known spin ice materials [38, 14, 39, 15]. These will theoretically lead to a further ordering into a long-range ordered phase approximately around $T_{c}\approx 0.07D_{nn}$, where $D_{nn}$ is the nearest-neighbor dipole-dipole interaction [14, 15]. It remains to be investigated how these long-range dipolar interactions further change the ground-state behavior also of the anisotropic model.

### v2

@@ V.1 Monte Carlo sampling
We collect states by Monte Carlo sampling a system of size $L=10$ (corresponding to $N_{\text{spins}}=16000$ spins) and compute the neutron scattering response. For $T\rightarrow 0$ states can be directly computed by sampling within the ground state manifold. To this end, sampling algorithms that allow us to generate all possible ground states of an anisotropic phase are employed. These sampling algorithms have a transition probability of $1$ when starting from a state within the ground state manifold and are described in detail in App. F.
@@
V.2 Self-Consisteng Gaussian Approximation
@@
Here, $L$ is the linear system size, and we insert the Fourier transform of our interaction Hamiltonian, specified in App. D. The Lagrange multiplier $\lambda$ can be interpreted as the thermal occupation value of the lowest energy mode [69] and will drop to zero if there is a phase transition in SCGA [69, 67]. An analysis of the Lagrange multiplier in the different $\delta$-regimes can be found in App. D and is consistent with the findings of section IV.
The scattering result in the SF channel in the SCGA is computed by inserting the correlator Eq. (8) into Eq. (6). To resolve this at different temperatures, we first need to solve for $\lambda$ in (9) at the necessary inverse temperature $\beta$. We can also compare the result from the SCGA to numerical data, where we sample ground states for each system according to the ground state rule. In the SCGA, we compute the correlations at a temperature well below the crossover ($T\ll 1J$) to accurately reproduce correlations close to or at the ground state.
@@ V.3 Results
The normalized SF channel of the neutron scattering cross section for the six different phases at or close to the ground state is shown in Fig. 4. The intensity is normalized for each system, $(S_{\text{SF}}(\mathbf{q})-\text{min}(S_{\text{SF}}(\mathbf{q})))/(\text{max}(S_{\text{SF}}(\mathbf{q}))-\text{min}(S_{\text{SF}}(\mathbf{q})))$, such that the signatures will lie between 0 and $1$. We can see that for any limit of $\delta_{A},\delta_{B}$, the spin structure factor computed numerically shows no significant difference from the result obtained by the SCGA when the temperature is set close to the transition temperature. It is important to note that the signatures on the antidiagonal in the [hl0] cut are due to the neutron polarization being almost parallel to the scattering wave vector and not a signature of the underlying spin configuration.
Figure 4: Normalized spin structure factor in Spin-Flip (SF) channel. Columns show the same system, and rows have the same cut in the Brillouin zone. Columns : 1 Isotropic, 2 Plane Ice, 3 Omni-Plane, 4 Line-Order, 5 Plane-Paramagnetic, 6 Ferromagnetic. Rows: 1 [hhl], 2 [hl0], 3 [h0l]. Every panel shows Monte Carlo (MC) data on the top and self-consistent Gaussian approximation (SCGA) on the bottom. The Monte Carlo data is obtained by sampling in the degenerate ground-state sector for systems with $L_{x}=L_{y}=L_{z}=10$ fcc unit cells, corresponding to $N_{\text{spin}}=16000$. The intensities are normalized across a single system between 0 and $1$ in a.u..
@@
In the planar ice phase, the pinch points with bow-tie patterns are no longer present in the [hhl] and [h0l] plane. Only in the [hl0] cut do pinch points still exist with a bow-tie pattern. This is understood from the cut being exactly parallel to the ice plane. Here, in the plane, correlations are still of dipolar form, with $C_{\alpha\beta}(\mathbf{r})\propto\frac{1}{r^{4}}(r^{2}\delta_{\alpha\beta}-3r_{\alpha}r_{\beta})$, for $r$ in the [hl0] plane. This will lead to a structure factor of the form $S_{\alpha\beta}\propto(\delta_{\alpha\beta}-\frac{k_{\alpha}k_{\beta}}{k^{2}})$ [75] in the plane, due to the square ice constraint in the ground state. The location of the pinch points in square ice are located at the high symmetry points of the square lattice, $(\pm 1,0)$, $(0,\pm 1)$, and $(\pm 1,\pm 1)$. This translates to pinch points in the [hl0] cut at $(\pm 100)$, $(0\pm 10)$, and $(\pm 1\pm 10)$ and symmetry equivalent points in the extended Brilloin zone. Since single planes are not correlated, other cuts in the lattice do not show this behavior.
@@
In the omni-plane case, spins order in particular patterns within the different tetrahedra-planes of the lattice. A single $A$ tetrahedra has a local magnetization vector that is either parallel or antiparallel to the $z$ axis. These tetrahedra will align predominantly antiferromagnetically along planes perpendicular to the $z$ direction and predominantly ferromagnetically for planes perpendicular to the $x$ and $y$ directions. A detailed description and graphical representation of this can be found in App. B.
@@
In a unit cell, there are six possible spin configurations. Two of which are fully ferromagnetic, leading to sharp Bragg peaks in the Brillouin zone center at $(000)$. The other four have all the $A$ tetrahedra in either the $x,z$ or the $y,z$-plane ferromagnetically aligned, while in all other planes, the $A$ tetrahedra are antiferromagnetically aligned. For these four states, however, two adjacent ferromagnetic layers have opposite polarizations. This is necessary to get the antiferromagnetic alignment in the other layers. When we now project the two ferromagnetic layers of a unit cell down onto the plane, the two oppositely polarized planes form an antiferromagnetic square lattice. For this order, characteristic Bragg peaks at the four corners of the square lattice can be observed. On the other hand, projecting the antiferromagnetic layers of a unit cell onto the plane gives a square lattice with stripe order. For a square lattice with horizontal or vertical stripe order, Bragg peaks at $(0,\pm 1)$ or $(\pm 1,0)$ appear.
@@
Any plane flip in the system will introduce line defects in the square lattice and thereby broaden the sharp Bragg peaks along the flipping direction. This explains the square grid structure in the [h0l] plane; the AFM square lattice peaks are broadened by introducing line defects. In the [hl0] plane, on the other hand, the stripe order peaks at $(\pm 1\pm 10)$ are broadened due to the introduction of line defects and form the cross-patterns. A detailed explanation with schematic figures can be found in App. B.
@@
In the plane-paramagnetic case, spins form decoupled planes, where within planes, the local magnetization vector of the $B$ tetrahedra align antiferromagnetically. This order in the $x,y$-planes is visible as Bragg peaks in the [hl0] cut, where the $B$ tetrahedra form a square lattice. In a square lattice with AFM order, we expect Bragg-peaks to form at $(\pm 1,\pm 1)$; however, the square lattice formed by the $B$ tetrahedra of a unit cell is rotated by $\pi/2$ compared to the $x,y$-projection of the fcc unit cell. This leads to the peaks at the expected locations of $(\pm 100)$ and $(0\pm 10)$ as well as symmetry equivalent positions in the extended Brillouin zone. The uncorrelated planes can be inferred by the vertical lines in the [h0l] cut.
@@
Figure 5: SCGA spin structure factor in Spin-Flip (SF) channel at different temperatures. While for high temperatures ($T\gg 1$) the signatures for the isotropic (left) and anisotropic (right) systems agree, when the temperature is lowered (top to bottom), the features look significantly different. The line-order regime is evaluated with $\delta_{A}=\delta_{B}=0.05$ and the [hhl] cut in the Brillouin zone is shown. The scale is chosen uniformly across all temperatures and both systems.
While these signatures will be visible at low temperatures, very close to the ground state, remnants of the change in correlations will also be visible at higher temperatures. We show how the neutron scattering response resulting from the SCGA changes between isotropic spin ice and the anisotropic systems for decreasing temperatures with the example of the line-order case in Fig. 5. Here, the colorscale is chosen uniformly across all temperatures and both systems for better comparison. We can observe that at high temperatures, for $T\gg 1J$, the scattering data will be indistinguishable between the isotropic and the anisotropic cases. However, significant differences can be observed below the spin-ice crossover, where the anisotropic systems gradually enter their respective ground-state manifold and the aforementioned signatures become visible. Measuring the structure factor as a function of temperature thus directly unveils the anisotropic couplings of the model Eq. (1).
@@ VI Summary and Outlook
In this work, we explored how anisotropic perturbations in the nearest-neighbor spin ice Hamiltonian affect the ground state structure and equilibrium properties of the spin system on the breathing pyrochlore lattice. By tuning the strain parameter, $\delta_{A/B}$, independently on the two tetrahedral sublattices, we demonstrated that the ground state degeneracy exhibits a sign-dependent change, leading to an additional five distinct phases. In all cases, the degeneracy is reduced; strikingly, in only one of the five anisotropic phases, the degeneracy grows exponentially with the full system volume, while in the other scenarios, the ground state degeneracy scales with a lower-dimensional cut of the system. Additionally, we analyzed the mobility of excitations on top of the ground state. We found that any negative $\delta$ rendered excitations immobile, while a positive $\delta$ made their mobility restricted to lower-dimensional submannifolds. These behaviors mirror key features of fracton systems, where excitations are immobile or exhibit constrained dynamics due to conservation laws or energetic barriers. In our case, the excitations are energetically restricted to planes or lines, or are completely frozen, making them fracton-like in nature.
@@
We also investigated the equilibrium properties of the system via the specific heat and entropy. Our results show that the anisotropic perturbation will either lead to a phase transition into a symmetry-broken phase at a critical temperature (omni-plane, plane paramagnetic, and ferromagnetic phase), a phase transition at zero temperature (line-order), or a crossover into a phase with decoupled planes (planar ice). In all cases but one, the planar ice phase, the residual entropy density approaches zero for increasing system sizes. The static spin structure factors of the anisotropic systems display significant deviations from the conventional spin ice case, showing only dipolar correlations in the planar ice phase.
Our analysis introduces a new perspective on frustrated magnetic systems, particularly in the context of strain engineering. While strain effects in spin ice systems have been previously studied [20, 19, 30, 17], introducing sublattice-dependent anisotropy enriches this field by opening a qualitatively different route to dimensional reduction. A well-known example of such reduction is the kagome ice state, realized under an external magnetic field along the global $[111]$ direction [76, 77, 78, 79, 80]. In that case, spins aligned with the field are pinned while the remaining spins fluctuate within decoupled two-dimensional kagome planes. By contrast, in our model, the reduction arises not from field-induced polarization but from frustration-lifting via sublattice-dependent bond anisotropy, providing an alternative mechanism for emergent lower-dimensional dynamics in spin ice.
@@
Looking forward, several intriguing questions arise from our findings. Since the ground-state degeneracy is significantly reduced in some cases, an important question concerns the mechanisms by which equilibrium is achieved. Furthermore, the timescales on which equilibrium is approached are to be investigated. Additionally, we analyzed the mobility of excitations on top of the ground states, which made a connection to fractons apparent [44, 45, 46, 47, 48, 24, 25]. This raises the question of how these energetic constraints and subdimensional mobility influence the relaxation dynamics of the system.
Finally, in this study, we restricted our analysis to the nearest-neighbor Ising Hamiltonian on the pyrochlore lattice, neglecting the long-range dipolar interactions that are intrinsic to real spin ice materials [51, 14, 52, 15]. These long-range interactions are known to induce a transition into a long-range ordered state at temperatures around $T_{c}\approx 0.07D_{nn}$, where $D_{nn}$ is the nearest-neighbor dipolar interaction strength [14, 15]. The dipolar spin ice model exhibits a set of twelve symmetry-related ground states, each characterized by a sixteen-site magnetic unit cell.
While our anisotropic model neglects these long-range terms, it still selects a smaller subset of states from the spin ice manifold, effectively reducing the degeneracy. Interestingly, we find that the ground-state construction rules of the anisotropic model generate states that, in most cases, are also ground states of the dipolar model; except in the ferromagnetic regime ($\delta_{A},\delta_{B}<0$), where full alignment conflicts with the long-range dipolar correlations. This indicates that there exists a class of configurations that are simultaneously ground states of both the classical spin ice model with long-range interactions and the anisotropic nearest-neighbor model with bond-dependent couplings.
This overlap suggests that introducing sublattice-resolved strain, even in the presence of dipolar interactions, could help lift the twelve-fold degeneracy of the dipolar model and potentially enhance the ordering temperature, providing a promising route toward controlled stabilization of specific spin ice ground states.
@@ Appendix A Experimental considerations
We now discuss possible realizations of the presented toy model. The system Hamiltonian (3) is based on the classical spin ice model. Due to large crystal field anisotropies, the spins are constrained to point along the local $\langle 111\rangle$ directions connecting each corner point of the tetrahedron to its center. This reduces the full spin Hamiltonian to an Ising-like interaction between the spins on each tetrahedron [2, 1]. Importantly, we consider a minimal model containing only the nearest-neighbor exchange interaction. In rare-earth spin ice compounds such as Dy2Ti2O7 and Ho2Ti2O7, magnetic interactions include both a short-range antiferromagnetic exchange $J_{\text{nn}}$ and a long-range ferromagnetic dipolar term $D_{\text{nn}}$. These combine into an effective nearest-neighbor coupling $J_{\text{eff}}=J_{\text{nn}}+D_{\text{nn}}$, which stabilizes the spin ice regime at intermediate temperatures. However, in our theoretical approach, we adopt a minimal description by retaining only the nearest-neighbor interactions and neglecting long-range dipolar terms beyond first neighbors. This approximation captures the correct thermodynamics at higher temperatures but omits the low-temperature ordering transition seen in full dipolar spin ice, which occurs around $T_{c}\approx 0.077D_{\text{nn}}$ [14]. Despite this simplification, the minimal model remains a valuable tool for exploring the role of anisotropy and frustration in modifying the spin ice ground-state manifold.